OpenStudy (anonymous):

how would you do ln6-ln2 without a calculator?

OpenStudy (amistre64):

what do you mean by "do"?

OpenStudy (mertsj):

That would be ln3

OpenStudy (anonymous):

*find the value

OpenStudy (amistre64):

given logs of the same base; they act like exponents. Exponents subtract when you divide like bases: \[\frac{b^n}{b^m}=b^{n-m}\] \[log_B(n)-log_B(m)=log_B(\frac{n}{m})\]

OpenStudy (amistre64):

ln6 - ln2 = ln 6/2 = ln3

OpenStudy (anonymous):

so you just move it over? thanks this really helps!

OpenStudy (amistre64):

yep, it "looks like" you move it over, but that is just the result of the math behind it :)

OpenStudy (anonymous):

so on this next one, 2ln6. It's completely different?

OpenStudy (amistre64):

correct, that is a different set up that maths up differently.

OpenStudy (amistre64):

the product of a constant and a log is equal to the log of the argument raised to the constant as a power: \[c\ log_B(N)=log_B(N^c)\]

OpenStudy (amistre64):

2 ln(6) = ln(6^2) = ln(36)

OpenStudy (anonymous):

Hmm, I see. I sort of understand what you're saying. 6ln2 would be the same, just flipped. I'm guessing.

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

the relationship between logs and exponents is: \[log_B(n)=m\ <->\ B^m = n\]

OpenStudy (amistre64):

they are inverses of each other and therefore share similar properties

OpenStudy (anonymous):

mk, ln6+ln2, ln6/2?

OpenStudy (amistre64):

addition is the opposite of subtraction; you notice that we divide when they subtract ... when they add, what do you think would be the opposite result?

OpenStudy (anonymous):

multiplication c:

OpenStudy (amistre64):

yep :) ln(6) + ln(2) = ln(6*2)

OpenStudy (anonymous):

that's it? ln(12)

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

yayayay

OpenStudy (anonymous):

log1/3 would beeee a subtraction then?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

log1-log3.

OpenStudy (anonymous):

Lol I think.

OpenStudy (amistre64):

log1-log3 = log1/3 and vice versa

OpenStudy (anonymous):

Math genius.

OpenStudy (anonymous):

l

OpenStudy (anonymous):

log48?